# RefAgent: A Multi-agent LLM-based Framework for Automatic Software Refactoring

### v1

@@ 4.3.1. Experimental details
We employ RefactoringMiner 3.0 (9136878; 8453111), which is a tool designed only to detect and classify refactoring operations between version histories or commits, to examine the refactorings applied by RefAgent. RefactoringMiner represents the state of the art in refactoring detection, supporting up to 59 different types of refactorings. The evaluation of RefactoringMiner on this dataset, performed in previous work (10.1109/TSE.2023.3326775), shows an overall precision of 99.7% and a recall of 94.2%, confirming it as an accurate refactoring detection tool (10.1145/3180155.3180206). The process is as follows: for each project $p$ at release version $v$, we create a fork. After RefAgent refactors a class, we commit the modified class to the fork.
@@ 4.3.2. Relevance to Search-based Refactoring Approaches and Developers
Previous work has shown that search-based refactoring approaches excel in identifying refactoring opportunities by using optimization algorithms that explore the search space of possible refactorings to maximize software quality (ALDALLAL2015231). Thus, we compare the refactoring opportunities identified by RefAgent with those made by RefGen, a search-based refactoring tool that implements efficient refactoring scheduling based on partial order reduction (MORALES201825). RefGen was initially developed as an Eclipse plug-in designed to suggest sequences of refactorings to improve the design quality of software systems by addressing anti-patterns and code smells detected across different classes. The generated refactoring sequence is ordered to maximize design quality improvement while avoiding conflicts among the suggested refactorings.
@@ 5.1.2. Approach
The gain in each metric is assessed using the Improvement Rate (IR), following the approach used in previous studies (cordeiro2024empiricalstudycoderefactoring). It is calculated as follows:
@@
To assess the statistical differences in improvements for code smells and unit tests relative to the baseline across all projects, we use the Wilcoxon signed-rank test (10.1145/1985793.1985795), as the measurements being compared are paired. A significance threshold of p <0.05<0.05 is used to determine statistical significance.
@@ 5.2.2. Approach
To evaluate RefAgent’s ability in identifying refactoring opportunities, we compare its refactoring selections with those made by developers and RefGen (MORALES201825), a search-based refactoring algorithm that was shown to effectively identify refactoring opportunities. The objective is to determine whether RefAgent applies refactorings in the same class, method, and location as developers and RefGen.
@@ 5.2.3. Findings
RefAgent
GPT-4o
RefAgent
Starcoder
RefAgent
Deepseek-coder
RefGen
@@ 5.3.1. Motivation
While multi-agent architectures offer modularity and the ability to decompose complex tasks, they also introduce design and orchestration overhead. Without a clear understanding of the role of each component, it becomes difficult to justify this complexity or optimize the framework. Prior work in software engineering and LLM-based systems highlights the importance of component-level analysis and ablation studies to quantify the effectiveness of individual elements (kang2023multimodal).
@@
To assess the performance of RefAgent, we compare its refactored outputs against single-agent approaches, which are based on previous work (cordeiro2024empiricalstudycoderefactoring), in terms of software quality improvements. The comparison focuses on key metrics used in RQ1, RQ2 ( unit test Pass Rate, compilation success rate, Smell Reduction Rate (SRR), and QMOOD metrics). For the single-agent baseline, we evaluate the effectiveness of an LLM-based refactoring agent given the same context as RefAgent to ensure a fair comparison. Similarly to previous work, we use for the unit test pass rate evaluation the pass @k metric defined as follows:
@@ 5.3.3. Findings
DeepSeek
coder
Pass@3 58 61 63 65.2 49.2 51.6
@@ 6. Discussion
To assess the reliability of Evosuite, we randomly select a statistically significant sample, averaging 374 classes per project from each of the eight open-source projects (95% confidence, 5% margin of error). For each class, we manually review all generated test cases to determine whether they were True Positives (TP), i.e., semantically correct and passing, or False Positives (FP), i.e., incorrect but still passing. As shown in Figure 9, an average of 91% of tests were TP, closely aligning with EvoSuite’s reported accuracy and prior studies (cordeiro2024empiricalstudycoderefactoring; shamshiri2015evosuite; 10.1145/3526072.3527526). The average FP rate was only 1.2%, confirming the low risk of misleading results. These findings support the reliability of EvoSuite within RefAgent. While FPs are a known limitation in automated test generation, their minimal presence suggests a limited impact on decision-making. Future work may incorporate mutation testing or enhanced test oracles to better validate behavioral correctness. Additionally, leveraging LLM-as-a-judge approaches can help assess semantic soundness beyond pass/fail outcomes (chen2023codellamaeval).
@@
For refactoring detection, as shown in Figure 9, RefactoringMiner yielded a TP rate of 98.24% and a FP rate of just 0.12%, confirming its reliability of the tool being aligned with the results reported by the authors of RefactoringMiner (9136878) However, a few false positives, typically benign cases like formatting changes being misclassified as refactorings, may introduce minor noise.
For code smells, as shown in Figure 9, DesigniteJava reported a TP rate of 93.6% (implementation smells) and 92.9% (design smells), with corresponding FP rates of 1.64% and 1.73% respectively. These results align with prior works (Sharma2024). While generally low, these false positives can slightly overestimate the number of smells detected or removed, particularly in borderline cases.
These findings highlight a broader challenge in automated evaluations: the trade-off between scalability and semantic precision. Since RefAgent operates as a fully automated framework, developers are not required to run static analysis tools before or after refactoring. Instead, we employ these tools solely for empirical evaluation, to quantitatively assess the improvements in code quality (e.g., reduction of code smells, refactoring types). The developer is not involved in this validation loop. In practice, RefAgent autonomously performs refactoring, testing, and validation using agent reasoning and tool-calling when needed. Therefore, the manual effort required from developers is significantly reduced. However, this paves the way for future work to introduce a human-in-the-loop review mechanism (takerngsaksiri2025humanintheloopsoftwaredevelopmentagents) that interacts with agents, which can further ensure reliability. An example is an agent-human collaboration mechanism in which it can request input at specific conversation rounds based on the configuration. The default user proxy agent should enable customizable involvement, allowing users to define how often and under what conditions human input is requested, including the option to skip providing input. These strategies offer a promising path toward balancing automation with trustworthiness in evaluating intelligent software engineering systems.
@@ 7. Threats to Validity
External Validity. In this study, we focused exclusively on datasets from Apache projects, which may limit the generalizability of our findings across different software development environments. However, this choice ensures consistency with related work (cordeiro2024empiricalstudycoderefactoring), enabling direct comparison and enhancing the validity and comparability of our results. Future research should explore more diverse datasets to further validate and broaden the applicability of our findings.

### v2

@@ 4.3.1. Experimental details
We employ RefactoringMiner 3.0 (Tsantalis et al., 2022, 2018b), which is a tool designed only to detect and classify refactoring operations between version histories or commits, to examine the refactorings applied by RefAgent. RefactoringMiner represents the state of the art in refactoring detection, supporting up to 59 different types of refactorings. The evaluation of RefactoringMiner on this dataset, performed in previous work (Noei et al., 2023), shows an overall precision of 99.7% and a recall of 94.2%, confirming it as an accurate refactoring detection tool (Tsantalis et al., 2018a). The process is as follows: for each project $p$ at release version $v$, we create a fork. After RefAgent refactors a class, we commit the modified class to the fork.
@@ 4.3.2. Relevance to Search-based Refactoring Approaches and Developers
Previous work has shown that search-based refactoring approaches excel in identifying refactoring opportunities by using optimization algorithms that explore the search space of possible refactorings to maximize software quality (Al Dallal, 2015). Thus, we compare the refactoring opportunities identified by RefAgent with those made by RefGen, a search-based refactoring tool that implements efficient refactoring scheduling based on partial order reduction (Morales et al., 2018). RefGen was initially developed as an Eclipse plug-in designed to suggest sequences of refactorings to improve the design quality of software systems by addressing anti-patterns and code smells detected across different classes. The generated refactoring sequence is ordered to maximize design quality improvement while avoiding conflicts among the suggested refactorings.
@@ 5.1.2. Approach
The gain in each metric is assessed using the Improvement Rate (IR), following the approach used in previous studies (Cordeiro et al., 2024). It is calculated as follows:
@@
To assess the statistical differences in improvements for code smells and unit tests relative to the baseline across all projects, we use the Wilcoxon signed-rank test (Arcuri and Briand, 2011), as the measurements being compared are paired. A significance threshold of p <0.05<0.05 is used to determine statistical significance.
@@ 5.2.2. Approach
To evaluate RefAgent’s ability in identifying refactoring opportunities, we compare its refactoring selections with those made by developers and RefGen (Morales et al., 2018), a search-based refactoring algorithm that was shown to effectively identify refactoring opportunities. The objective is to determine whether RefAgent applies refactorings in the same class, method, and location as developers and RefGen.
@@ 5.3.1. Motivation
While multi-agent architectures offer modularity and the ability to decompose complex tasks, they also introduce design and orchestration overhead. Without a clear understanding of the role of each component, it becomes difficult to justify this complexity or optimize the framework. Prior work in software engineering and LLM-based systems highlights the importance of component-level analysis and ablation studies to quantify the effectiveness of individual elements (Nunez et al., 2024a).
@@
To assess the performance of RefAgent, we compare its refactored outputs against single-agent approaches, which are based on previous work (Cordeiro et al., 2024), in terms of software quality improvements. The comparison focuses on key metrics used in RQ1, RQ2 ( unit test Pass Rate, compilation success rate, Smell Reduction Rate (SRR), and QMOOD metrics). For the single-agent baseline, we evaluate the effectiveness of an LLM-based refactoring agent given the same context as RefAgent to ensure a fair comparison. Similarly to previous work, we use for the unit test pass rate evaluation the pass @k metric defined as follows:
@@ 5.3.3. Findings
DeepSeek coder Pass@3 58 61 63 65.2 49.2 51.6
@@ 6. Discussion
To assess the reliability of Evosuite, we randomly select a statistically significant sample, averaging 374 classes per project from each of the eight open-source projects (95% confidence, 5% margin of error). For each class, we manually review all generated test cases to determine whether they were True Positives (TP), i.e., semantically correct and passing, or False Positives (FP), i.e., incorrect but still passing. As shown in Figure 9, an average of 91% of tests were TP, closely aligning with EvoSuite’s reported accuracy and prior studies (Cordeiro et al., 2024; Shamshiri et al., 2015; Schweikl et al., 2023). The average FP rate was only 1.2%, confirming the low risk of misleading results. These findings support the reliability of EvoSuite within RefAgent. While FPs are a known limitation in automated test generation, their minimal presence suggests a limited impact on decision-making. Future work may incorporate mutation testing or enhanced test oracles to better validate behavioral correctness. Additionally, leveraging LLM-as-a-judge approaches can help assess semantic soundness beyond pass/fail outcomes (Gu et al., 2025).
@@
For refactoring detection, as shown in Figure 9, RefactoringMiner yielded a TP rate of 98.24% and a FP rate of just 0.12%, confirming its reliability of the tool being aligned with the results reported by the authors of RefactoringMiner (Tsantalis et al., 2022) However, a few false positives, typically benign cases like formatting changes being misclassified as refactorings, may introduce minor noise.
For code smells, as shown in Figure 9, DesigniteJava reported a TP rate of 93.6% (implementation smells) and 92.9% (design smells), with corresponding FP rates of 1.64% and 1.73% respectively. These results align with prior works (Sharma, 2024). While generally low, these false positives can slightly overestimate the number of smells detected or removed, particularly in borderline cases.
These findings highlight a broader challenge in automated evaluations: the trade-off between scalability and semantic precision. Since RefAgent operates as a fully automated framework, developers are not required to run static analysis tools before or after refactoring. Instead, we employ these tools solely for empirical evaluation, to quantitatively assess the improvements in code quality (e.g., reduction of code smells, refactoring types). The developer is not involved in this validation loop. In practice, RefAgent autonomously performs refactoring, testing, and validation using agent reasoning and tool-calling when needed. Therefore, the manual effort required from developers is significantly reduced. However, this paves the way for future work to introduce a human-in-the-loop review mechanism (Takerngsaksiri et al., 2025) that interacts with agents, which can further ensure reliability. An example is an agent-human collaboration mechanism in which it can request input at specific conversation rounds based on the configuration. The default user proxy agent should enable customizable involvement, allowing users to define how often and under what conditions human input is requested, including the option to skip providing input. These strategies offer a promising path toward balancing automation with trustworthiness in evaluating intelligent software engineering systems.
@@ 7. Threats to Validity
External Validity. In this study, we focused exclusively on datasets from Apache projects, which may limit the generalizability of our findings across different software development environments. However, this choice ensures consistency with related work (Cordeiro et al., 2024), enabling direct comparison and enhancing the validity and comparability of our results. Future research should explore more diverse datasets to further validate and broaden the applicability of our findings.